OpenStudy (anonymous):

find the sum of the first six terms of the geometric sequence for which a2 = .7 and a3= .49

OpenStudy (ranga):

Common ratio \(\Large r = \frac{a_3}{a_2} = ?\)

OpenStudy (anonymous):

.7

OpenStudy (ranga):

Yes. Find the first term a1 from 0.7 = a2 / a1 = .7 / a1

OpenStudy (anonymous):

0?

OpenStudy (ranga):

0.7 = 0.7 / a1. a1 = 0.7 / 0.7 = 1. Plug it into the formula for the sum of a geometric series. a1 = 1; r = 0.7; n = 6

OpenStudy (anonymous):

I got -5/.3 is that correct? @ranga

OpenStudy (ranga):

Sum = \[ a_1 * \frac{1-r^n}{1-r} = 1 * \frac{1-0.7^6}{1-0.7} = ?\]

OpenStudy (anonymous):

2.94117

OpenStudy (anonymous):

thank you, you are a life saver!!

OpenStudy (ranga):

Yes.